OpenStudy (anonymous):

Solve for x. 4x(x-1)-5x(x)=3

OpenStudy (anonymous):

first distributive law \[4x^2-4x-5x^2=3\] then combine like terms \[-x^2-4x=3\] the a quadratic equation \[x^2+4x+3=0\] from here you can factor and solve for x

OpenStudy (anonymous):

-1, -3